OpenStudy (anonymous):

cosx/1-tanx + sinx/1-cotx = sinx + cos x

OpenStudy (anonymous):

write tanx=sinx/cosx and cotx=cosx/sinx, then solve it

OpenStudy (anonymous):

there is an error in this question, this question can only be solved if one of the sign is +ve, you can't have them both negative, if you want them both negative then the answer would be -(sin(x) + cos(x))

OpenStudy (anonymous):

\[ \sin(x)/1-\cot(x) = \frac{ \sin(x) }{ \sin(x) - \cos(x) } \]

OpenStudy (anonymous):

similarly, cos(x)/1−tan(x)=cos(x)/[cos(x)−sin(x)]

OpenStudy (anonymous):

now look at the denominator, what happens when you multiply them both, consider the following

OpenStudy (anonymous):

isnt it supposed to be sin^2x and cos^2x as the numerators ?

OpenStudy (anonymous):

(m-k)(k-m) = \[-k ^{2}-m ^{2}-km +km\]

OpenStudy (anonymous):

yes hammy you are right but it is -1

OpenStudy (anonymous):

-sin^2(x)-cos^2(x)=-1

OpenStudy (anonymous):

you can figure that out just by looking at the signs

OpenStudy (anonymous):

if it was (m-k)(m+k) then you could get a +ve answer

OpenStudy (anonymous):

cos^2(x) at the numerator, i am just new to this editor, i;; get used to it

OpenStudy (anonymous):

alrighty then :) wellll i got as far as ( cos^2x/cosx-sinx) +(sin^2x /sinx- cosx0

OpenStudy (anonymous):

shoot the 0 sposed to be a closed bracket

OpenStudy (anonymous):

see i told u hehehe

OpenStudy (anonymous):

i need to know if this editor supports latex

OpenStudy (anonymous):

well i cant help you with that because i do not know sorry

OpenStudy (anonymous):

yes i think there is an error in the problem or probably the student didn't copy the question the question in the right way

OpenStudy (anonymous):

oh well thats that

OpenStudy (anonymous):

hey but i solved it and the ans came out alright

OpenStudy (anonymous):

u have to take the minus common from the denominator, then it will be \[\cos ^{2}x/(cosx-sinx) -\sin ^{2}x/(cosx-sinx)\] then just simple adding, u will get \[(\cos ^{2}x-\sin ^{2}x)/(cosx-sinx)\] then apply \[a ^{2}-b ^{2}= (a+b)(a-b)\] ie. \[(cosx+sinx)(cosx-sinx)/(cosx-sinx)\]= sinx+cosx